DarkSide-UK Cold test stand integration call




Zoom connection:


Outstanding/key items:

  • LTC6820 demo board availability is poor for steering module control boxes. Liverpool has 4 on order but delivery has slipped to DEC 20. Four multiplexer boards have been ordered as a backup.
  • Tile/vPDU testing status? Current estimate is that vPDUs will be made available for the cold test stands in March/April.
  • vPDU vacuum bagging
  • Discussion of cable testing at RHUL.
  • Planning for LV and HV(?) filtering
  • Are shielded cables required inside the detector?
    • 10:05 11:05
      Astrocent test stand updates 1h



      Astrocent will use CAEN VX2740 boards.

    • 10:05 11:05
      Edinburgh test stand updates 1h

      MIDAS is working. Will look at setup steps for other test stands.

    • 10:05 11:05
      Lancaster test stand updates 1h
    • 10:05 11:05
      Light source updates 1h

      Working on stiff design for connecting fiber to laser coupler.

      Task for Liverpool: Share diffuser positions and check with simulations if coverage is acceptable.

      Considering working towards a light source available late December if possible - TBD.

    • 10:05 11:05
      Liverpool test stand updates 1h

      Liverpool construction schedule:


    • 10:05 11:05
      RHUL test stand updates 1h
    • 10:05 11:05
      Signal cable and feedthrough plans 1h

      DB50 signal feedthroughs into CAEN VX2745 using custom HDMI adapter PCB for pin headers on front panel.

      DB50 power feedthroughs into either CAEN or desktop Keithley/Keysight power supplies.

      Grounding and cable shielding.